Not logged in
Tue Aug 8 20:48:06 2017, gautam, Update, SUS, ITMX stuck
Thu Aug 10 09:31:57 2017, Steve, Update, SUS, ITMX sensor voltage
Tue Aug 8 20:48:06 2017
Reply to this:
Somewhere between CDS model restarts and the IFO venting, ITMX got stuck
I shook it loose using the usual bias slider technique. It appears to be free now, I was able to lock the green beam on a TEM00 mode without touching the green input pointing. The ITMX Oplev spot has also returned to within its MEDM display bounds.